# Multi-Grained Query-Guided Set Prediction Network for Grounded Multimodal Named Entity Recognition

### v2

@@ Conclusion
In this paper, we propose a novel method named MQSPN to model appropriate intra-entity and inter-entity relationships for the GMNER task. To achieve this, we propose a learnable Multi-grained Queries Set (MQS) to adaptively learn explicit intra-entity connections. Besides, we reformulate GMNER as a multimodal set prediction (MSP) to model inter-entity relationships from an optimal matching view. Experimental results demonstrate that MQSPN achieves state-of-the-art performance on GMNER and its 2 subtasks across two Twitter benchmarks. In the future, the extension of our MQSPN to the vision-language pre-training model or knowledge-enhanced framework in few-shot or zero-shot scenarios will be one of our directions.
@@
Actually, MQSPN also achieves higher performance upper bound by employing stronger foundation models as its components. We present the performance of MQSPN with stronger foundation models in GMNER and its two subtasks (i.e. MNER and EEG) across Twitter-GMNER and Twitter-FMNERG datasets, as shown in Table 4. Specifically, we additionally choose xlm-roberta-large [^5] https://hf-mirror.com/FacebookAI/xlm-roberta-large (Roberta-large) for text encoder, clip-vit-large-patch14666https://hf-mirror.com/openai/clip-vit-large-patch14 (ViT-L/14) for vision encoder, and Faster-RCNN [^7] https://github.com/open-mmlab/mmdetection for region proposal network. Different combinations of components are considered to be different variants of MQSPN.
@@
The GMNER prediction is composed of entity span, type, and visual region. Following previous works(Yu et al. 2023), The correctness of each prediction is computed as follows:
@@
The first group first uses any previous state-of-the-art MNER method and object detector, i.e., VinVL (Zhang et al. 2021c) or Faster R-CNN (Girshick 2015), to pre-extract entity span-type pairs and type-related candidate regions. Entity-aware Visual Grounding (EVG) module (Yu et al. 2023) is then used to predict the matching relationship between entities and regions. The following MNER models are utilized as strong baselines: (1) GVATT (Lu et al. 2018b) adopts a visual attention mechanism based on BiLSTM-CRF to extract multimodal entities. (2) UMT (Yu et al. 2020) proposes a multimodal transformer to capture the cross-modality semantics. (3) UMGF (Zhang et al. 2021a) solves text-image integration with a multimodal graph fusion mechanism. (4) ITA (Wang et al. 2022b) utilizes the image-text translation and object tags to explicitly align visual and textual features. (5) BARTMNER (Yu et al. 2023) expands generative model BART with cross-modal transformer layer.
The second group includes several unified GMNER approaches. (1) H-Index (Yu et al. 2023) uses a multimodal BART model with the pointer mechanism to formulate the GMNER task as a sequence generation. (2) TIGER (Wang et al. 2023) is T5-based generative model which converts all span-type-region triples into target paraphrase sequences. (3) MNER-QG (Jia et al. 2023) is a unified MRC framework that combines textual entity extraction and grounding with multi-task joint learning. Please note that we replace the weakly-supervised query grounding in MNER-QG with a fully-supervised mode.

### v3

@@ Conclusion
In this paper, we propose a novel method named MQSPN to model appropriate intra-entity and inter-entity relationships for the GMNER task. To achieve this, we propose a learnable Multi-grained Queries Set (MQS) to adaptively learn explicit intra-entity connections. Besides, we reformulate GMNER as a multimodal set prediction (MSP) to model inter-entity relationships from an optimal matching view. Experimental results demonstrate that MQSPN achieves state-of-the-art performance on GMNER and its 2 subtasks across two Twitter benchmarks.
Acknowledgements
This work is supported by the National Natural Science Foundation of China (U2001211, U22B2060, 62276279), National Natural Science Foundation Youth Basic Research Program (623B2100), Research Foundation of Science and Technology Plan Project of Guangzhou City (2023B01J0001, 2024B01W0004), Guangdong Basic and Applied Basic Research Foundation (2024B1515020032).
@@ Acknowledgements
Actually, MQSPN also achieves higher performance upper bound by employing stronger foundation models as its components. We present the performance of MQSPN with stronger foundation models in GMNER and its two subtasks (i.e. MNER and EEG) across Twitter-GMNER and Twitter-FMNERG datasets, as shown in Table 4. Specifically, we additionally choose xlm-roberta-large [^4] https://hf-mirror.com/FacebookAI/xlm-roberta-large (Roberta-large) for text encoder, clip-vit-large-patch14555https://hf-mirror.com/openai/clip-vit-large-patch14 (ViT-L/14) for vision encoder, and Faster-RCNN [^6] https://github.com/open-mmlab/mmdetection for region proposal network. Different combinations of components are considered to be different variants of MQSPN.
@@
The GMNER prediction is composed of entity span, type, and visual region. Following previous works(Yu et al. 2023a), The correctness of each prediction is computed as follows:
@@
The first group first uses any previous state-of-the-art MNER method and object detector, i.e., VinVL (Zhang et al. 2021c) or Faster R-CNN (Girshick 2015), to pre-extract entity span-type pairs and type-related candidate regions. Entity-aware Visual Grounding (EVG) module (Yu et al. 2023a) is then used to predict the matching relationship between entities and regions. The following MNER models are utilized as strong baselines: (1) GVATT (Lu et al. 2018b) adopts a visual attention mechanism based on BiLSTM-CRF to extract multimodal entities. (2) UMT (Yu et al. 2020) proposes a multimodal transformer to capture the cross-modality semantics. (3) UMGF (Zhang et al. 2021a) solves text-image integration with a multimodal graph fusion mechanism. (4) ITA (Wang et al. 2022) utilizes the image-text translation and object tags to explicitly align visual and textual features. (5) BARTMNER (Yu et al. 2023a) expands generative model BART with cross-modal transformer layer.
The second group includes several unified GMNER approaches. (1) H-Index (Yu et al. 2023a) uses a multimodal BART model with the pointer mechanism to formulate the GMNER task as a sequence generation. (2) TIGER (Wang et al. 2023) is T5-based generative model which converts all span-type-region triples into target paraphrase sequences. (3) MNER-QG (Jia et al. 2023) is a unified MRC framework that combines textual entity extraction and grounding with multi-task joint learning. Please note that since the source code of MNER-QG has not been released, we reproduced this method by following the same settings as MQSPN, based on its original paper. Specifically, we replaced the original YOLOv3 with VinVL and eliminated the complex cross-modal interaction module. Additionally, we replaced the weakly-supervised query grounding in MNER-QG with a fully-supervised mode.